# DCAT: Dual Cross-Attention Fusion for Disease Classification in Radiological Images with Uncertainty Estimation

### v1

@@ 2 Related work
Attention mechanisms have emerged as a powerful tool in computer vision. Attention mechanism can be broadly categorized into channel attention, spatial attention, and cross attention, each addressing specific aspects of feature representation. Channel attention mechanisms, such as Squeeze-and-Excitation (SE) networks  (Hu et al., 2018) and selective kernel attention  (Chen et al., 2017) allow models to recalibrate channel-wise feature responses, emphasizing important features adaptively. Spatial attention mechanisms, on the other hand, focus on identifying salient spatial regions within feature maps  (Woo et al., 2018; Dai et al., 2017). Techniques like CBAM  (Woo et al., 2018) and deformable convolution  (Dai et al., 2017) have significantly improved tasks requiring precise localization, such as object detection and segmentation. Self-attention mechanisms, popularized by Transformers  (Zhao et al., 2020), capture global dependencies by computing attention weights across all spatial positions  (Wang et al., 2023b). Variants like Swin Transformers  (Liu et al., 2021), Focal Transformers  (Yang et al., 2021), and HRFormer  (Yuan et al., 2021) have extended this capability to process multi-scale information effectively. However, these approaches often neglect adjusting channel dimensions, restricting their capacity to fully leverage channel-wise feature interactions. Cross-attention mechanisms have gained traction for their ability to model relationships between modalities or feature maps. Unlike self-attention, which operates within a single feature space, cross-attention enables exchange between distinct feature representations, enhancing the model’s ability to capture contextual dependencies. For instance,  (Schlemper et al., 2019) introduced an attention gate model for medical image analysis, which suppresses irrelevant regions while highlighting task-specific features. Similarly,  (Guan et al., 2020) proposed a three branch attention guided CNN for thorax disease identification, integrating global and local cues to improve performance. These studies underscore the potential of attention mechanisms in addressing challenges such as noise suppression and feature refinement. Despite advancements in attention mechanisms, several limitations persist, particularly in cross-attention multi-network feature fusion. While channel and spatial attention mechanisms have been studied extensively, their integration with cross-attention remains insufficiently explored. Finally, lack of uncertainty quantification in attention limits their reliability in critical applications like medical imaging.
@@ 3.1 Materials
In our image classification task, we evaluated the proposed DCAT model using four publicly available medical imaging datasets, Covid-19 chest X-ray  (Chowdhury et al., 2020), Tuberculosis chest X-ray  (Rahman et al., 2020), Pneumonia Chest X-ray  (Kermany, 2018), OCT  (Kermany, 2018) images. Figure 1(a) and (b) shows some randomly selected samples from four datasets and their class-wise distribution respectively. We randomly split the data into two main sets, where 80% of the data was used for training and the remaining 20% was used as test sets. Two important points to be considered regarding datasets in this work. In the Covid-19 chest X-ray dataset, the original datasets contain four classes (Covid-19, Lung-Opacity, Viral Pneumonia, and Normal). However, Viral Pneumonia, and Normal classes have similar samples as the Pneumonia chest X-ray datasets. To prevent data leakage and overlapping, we discard these two classes and have considered only the Covid-19 and the Lung Opacity classes from the dataset. The t-SNE plot in Figure 2 (a) and (b) gives the visualization and comparison of distributions across datasets to identify similarities (overlap) and differences.
Figure 3: Overview of the proposed DCAT fusion mechanism. The proposed mechanism starts with acquiring chest X-ray and OCT images. The proposed model comprised of local and global feature learning by EfficientNetB4 and ResNet34, cross-attention fusion, and predictive uncertainty estimation. And finally, we have uncertainty estimation quantifying the uncertainty of the model.
@@ 3.2 Proposed Methodology
This paper proposes a novel cross-attention fusion called DCAT model that integrates multi-scale feature extraction using EfficientNetB4 and ResNet34 backbones. As illustrated in Figure 3, the proposed architecture comprises four components: dual-branch feature extraction with EfficientNetB4 and ResNet34, hierarchical multi-scale and cross-attention fusion mechanism, refined features using enhanced CBAM attention mechanism, and classification and predictive uncertainty estimation.
EfficientNetB4, part of Efficient family introduced by Tan and Le (2019) in 2019 employs a compound scaling method that uniformly scales depth, width, and resolution to optimize performance with fewer parameters. This scaling enables the network to capture fine-grained details and high-level semantic information. ResNet34, part of the Residual Network (ResNet) family introduced by He et al. (2016), addresses the vanishing gradient problem through its residual learning framework, enabling the training of deep networks. The architecture extracts hierarchical features, with early layers capturing low-level details and deeper layers capturing high-level semantics. Its residual connections make it robust to over-fitting, even with limited training data. Combining these architectures, the proposed model leverages their complementary strengths in feature representations.
@@ 3.5 Cross Attention mechanism
Let $X^{eff}_{img}\in\mathbb{R}^{C\times H\times W}$ represent the feature map from EfficientNetB4, and $X^{r}_{img}\in\mathbb{R}^{C\times H\times W}$ represent the feature map from ResNet34. To enable bi-directional attention fusion, we compute $\text{Query}(Q),\text{Key}(K)$, and $\text{Value}(V)$ representations of both the networks. For EfficientNetB4, Q termed as $Q_{eff}$ is derived from ResNet34 features and are the transformed EfficientNetB4 features (shown in Figure 3). K and Q termed as $K_{eff}$ and $V_{eff}$ respectively are computed from EfficientNetB4 features.
@@ 3.6 Feature Refinement and Classification
Following the cross-attention fusion, we refine the feature maps using an enhanced version of the CBAM attention mechanism (Figure 4). By leveraging channel and spatial attention mechanism, our refined CBAM enables the model to attend to the channel and spatial axes, respectively.
@@ 4.1 Implementation details
We choose Pneumonia, Covid-19, Tuberculosis, and Normal for chest X-ray images belonging to three different datasets and CNV, DME, DRUSEN, and Normal for retinal OCT images as the categories of classification task. Our chest X-ray images are collected from public repositories. We implement the proposed mechanism and perform experiments using the PyTorch library. Training and evaluations have been accomplished using NVIDIA GeForce RTX 3060 Ti inbuilt GPU. The batch size for training is set to 32. The maximum epochs is set to 200 and the learning rate is initialized to $43e-$ 4 with a weight decay of $41e-$ 4. We choose the Adam as the optimizer for the stage of training. The whole training process takes 2h30min and the testing time for a per radiology image is approximately 0.015s.
@@ 4.1.1 Evaluation metrics
For a comprehensive overview and evaluation of performance of the proposed model, we employed standard performance metrics like accuracy, AUC is the area under the ROC curve (AUROC), Area Under the Precision-Recall (AUPR), precision, recall, specificity, F1-score, Matthews Correlation Coefficient (MCC) and Cohen’s Kappa. As our dataset is class imbalanced, we use the AUPR as a metric that evaluates classification performance. AUC, AUPR focuses on the positive class performance. MCC and Cohen’s Kappa are both metrics used to evaluate classification performance, but they have distinct differences in handling agreement and correlation. Since our datasets are class imbalanced, we considered MCC for evaluating the classifier performance. The Kappa coefficient, which is used to measure the classification accuracy of the model and can also be used for consistency testing.

### v2

@@ 2 Related work
Attention mechanisms have emerged as a powerful tool in computer vision. Attention mechanisms can be broadly categorized into channel attention, spatial attention, and cross attention, each addressing specific aspects of feature representation. Channel attention mechanisms, such as Squeeze-and-Excitation (SE) networks  (Hu et al., 2018) and selective kernel attention  (Chen et al., 2017) allow models to recalibrate channel-wise feature responses, emphasizing important features adaptively. Spatial attention mechanisms, on the other hand, focus on identifying salient spatial regions within feature maps  (Woo et al., 2018; Dai et al., 2017). Techniques like CBAM  (Woo et al., 2018) and deformable convolution  (Dai et al., 2017) have significantly improved tasks requiring precise localization, such as object detection and segmentation. Self-attention mechanisms, popularized by Transformers  (Zhao et al., 2020), capture global dependencies by computing attention weights across all spatial positions  (Wang et al., 2023b). Variants like Swin Transformers  (Liu et al., 2021), Focal Transformers  (Yang et al., 2021), and HRFormer  (Yuan et al., 2021) have extended this capability to process multi-scale information effectively. However, these approaches often neglect to adjust channel dimensions, restricting their capacity to fully leverage channel-wise feature interactions. Cross-attention mechanisms have gained traction for their ability to model relationships between modalities or feature maps. Unlike self-attention, which operates within a single feature space, cross-attention enables exchange between distinct feature representations, enhancing the model’s ability to capture contextual dependencies. For instance,  (Schlemper et al., 2019) introduced an attention gate model for medical image analysis, which suppresses irrelevant regions while highlighting task-specific features. Similarly,  (Guan et al., 2020) proposed a three-branch attention-guided CNN for thorax disease identification, integrating global and local cues to improve performance. These studies underscore the potential of attention mechanisms in addressing challenges such as noise suppression and feature refinement. Despite advancements in attention mechanisms, several limitations persist, particularly in cross-attention multi-network feature fusion. While channel and spatial attention mechanisms have been studied extensively, their integration with cross-attention remains insufficiently explored. Finally, the lack of uncertainty quantification in attention limits their reliability in critical applications like medical imaging.
@@ 3.1 Materials
In our image classification task, we evaluated the proposed DCAT model using four publicly available medical imaging datasets, Covid-19 chest X-ray  (Chowdhury et al., 2020), Tuberculosis chest X-ray  (Rahman et al., 2020), Pneumonia Chest X-ray  (Kermany, 2018), OCT  (Kermany, 2018) images. Figure 1(a) and (b) show some randomly selected samples from four datasets and their class-wise distribution respectively. We randomly split the data into two main sets, where 80% of the data was used for training and the remaining 20% was used as test sets. Two important points to be considered regarding datasets in this work. In the Covid-19 chest X-ray dataset, the original datasets contain four classes (Covid-19, Lung-Opacity, Viral Pneumonia, and Normal). However, Viral Pneumonia, and Normal classes have similar samples as the Pneumonia chest X-ray datasets. To prevent data leakage and overlapping, we discard these two classes and have considered only the Covid-19 and the Lung Opacity classes from the dataset. The t-SNE plot in Figure 2 (a) and (b) gives the visualization and comparison of distributions across datasets to identify similarities (overlap) and differences.
Figure 3: Overview of the proposed DCAT fusion mechanism. The proposed mechanism starts with acquiring chest X-ray and OCT images. The proposed model comprised of local and global feature learning by EfficientNetB4 and ResNet34, cross-attention fusion, and predictive uncertainty estimation. Finally, we have uncertainty estimation quantifying the uncertainty of the model.
@@ 3.2 Proposed Methodology
This paper proposes a novel cross-attention fusion called the DCAT model that integrates multi-scale feature extraction using EfficientNetB4 and ResNet34 backbones. As illustrated in Figure 3, the proposed architecture comprises four components: dual-branch feature extraction with EfficientNetB4 and ResNet34, hierarchical multi-scale and cross-attention fusion mechanism, refined features using enhanced CBAM attention mechanism, and classification and predictive uncertainty estimation.
EfficientNetB4, part of the Efficient family introduced by Tan and Le (2019) in 2019 employs a compound scaling method that uniformly scales depth, width, and resolution to optimize performance with fewer parameters. This scaling enables the network to capture fine-grained details and high-level semantic information. ResNet34, part of the Residual Network (ResNet) family introduced by He et al. (2016), addresses the vanishing gradient problem through its residual learning framework, enabling the training of deep networks. The architecture extracts hierarchical features, with early layers capturing low-level details and deeper layers capturing high-level semantics. Its residual connections make it robust to over-fitting, even with limited training data. Combining these architectures, the proposed model leverages their complementary strengths in feature representations.
@@ 3.5 Cross Attention mechanism
Let $X^{eff}_{img}\in\mathbb{R}^{C\times H\times W}$ represent the feature map from EfficientNetB4, and $X^{r}_{img}\in\mathbb{R}^{C\times H\times W}$ represent the feature map from ResNet34. To enable bi-directional attention fusion, we compute $\text{Query}(Q),\text{Key}(K)$, and $\text{Value}(V)$ representations of both networks. For EfficientNetB4, Q termed as $Q_{eff}$ is derived from ResNet34 features and are the transformed EfficientNetB4 features (shown in Figure 3). K and Q termed as $K_{eff}$ and $V_{eff}$ respectively are computed from EfficientNetB4 features.
@@ 3.6 Feature Refinement and Classification
Following the cross-attention fusion, we refine the feature maps using an enhanced version of the CBAM attention mechanism (Figure 4). By leveraging channel and spatial attention mechanisms, our refined CBAM enables the model to attend to the channel and spatial axes, respectively.
@@ 4.1 Implementation details
We choose Pneumonia, Covid-19, Tuberculosis, and Normal for chest X-ray images belonging to three different datasets and CNV, DME, DRUSEN, and Normal for retinal OCT images as the categories of classification task. Our chest X-ray images are collected from public repositories. We implement the proposed mechanism and perform experiments using the PyTorch library. Training and evaluations have been accomplished using NVIDIA GeForce RTX 3060 Ti inbuilt GPU. The batch size for training is set to 32. The maximum epochs is set to 200 and the learning rate is initialized to $43e-$ 4 with a weight decay of $41e-$ 4. We choose Adam as the optimizer for the training stage. The whole training process takes 2h30min and the testing time for a per radiology image is approximately 0.015s.
@@ 4.1.1 Evaluation metrics
For a comprehensive overview and evaluation of the performance of the proposed model, we employed standard performance metrics like accuracy, AUC is the area under the ROC curve (AUROC), Area Under the Precision-Recall (AUPR), precision, recall, specificity, F1-score, Matthews Correlation Coefficient (MCC) and Cohen’s Kappa. As our dataset is class-imbalanced, we use the AUPR as a metric that evaluates classification performance. AUC, AUPR focus on positive class performance. MCC and Cohen’s Kappa are both metrics used to evaluate classification performance, but they have distinct differences in handling agreement and correlation. Since our datasets are class imbalanced, we considered MCC for evaluating the classifier performance. The Kappa coefficient, is used to measure the classification accuracy of the model and can also be used for consistency testing.